# version 0.75

### v4

@@ 9.3.1 Linear Predictors
The simplest prediction rule is a linear function of the observations, inputs, or both. More precisely,
@@ 9.3.2 Linear Dynamical Predictors
The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $\mathbf{C}=I$ is the identity matrix. In this situation, the optimal prediction can be computed via the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$. It can be seen that this prediction rule is optimal for any zero-mean noise or perturbation model. However, this optimality is only true in the case of full observation of state. In the next chapter we consider optimal predictors for zero mean i.i.d noise in the more general partially observed case.
@@
Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of states and controls and gives the next system output accurately. We denote by $\prod^{\star}$ the set of all such predictors, that in the case of zero noise give a perfectly accurate prediction.
@@
With a decaying system, we can limit the memory of the predictor and yet maintain accuracy. Thus it makes sense to consider a more restricted class of predictors that have bounded memory, i.e.
@@
where we define all $\mathbf{u}$ vectors in negative times as zero. Denote by $\prod_{h}^{\star}$ the class of all such predictors.
Clearly for a large $h$ this class of predictors well approximates the true observations $\mathbf{y}_{t}$ if the system is activated without any noise. We formalize what this means next.
@@
$=\|\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t-1}\mathbf{x}_{0}-\sum_{i=1:h}CA^{i-1}B\mathbf{u}_{t-i}\|$
$=\|\sum_{i=h}^{T}CA^{i}B\mathbf{u}_{t-i-1}\|$
@@ 9.4 Efficient Learning of Linear Predictors
The previous section gives several natural classes of predictors as well some of the relationships between them. Notably, the class of linear dynamical predictors is optimal for certain scenarios and requires knowledge of only three matrices. However, the parametrization of this class is non-convex in the matrices $A,B,C$, which presents a computational challenge.
This challenge is circumvented by using the fact that the class $\prod^{L}$ $\varepsilon$-approximates the class $\prod^{\star}$. Further, the natural parametrization of $\prod^{L}$ is convex and thus can be learned using online convex optimization.
@@ 9.5 Summary
We began by formulating the problem in terms of partially observable LDS, highlighting the challenges posed by unknown system matrices and noisy observations. We discussed key approaches for online learning in dynamical systems, emphasizing regret minimization as a performance metric. The regret framework allows for direct comparisons between learning algorithms by evaluating their cumulative prediction error against the best fixed predictor in hindsight.
@@ 9.6 Bibliographic Remarks
The task of predicting future observations in an online environment is often called “filtering”. Resting at the crossroads between statistical estimation and dynamical systems, the theory of filtering had its early beginnings in the works of Norbert Wiener [Wie49].
For a comprehensive treatment of online learning in games and online convex optimization, see the texts [CBL06, Haz16].
In the statistical noise setting, a natural approach for learning is to perform system identification and then use the identified system for prediction. Such an approach was taken by [SMT+18, SF20, SR19]. The work of [GLS+20] extends identification-based techniques to adversarial noise and marginally stable systems.
The class $\prod^{*}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [AHL+18]. In that paper, it was shown how to learn partially observable, and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
@@ 9.7 Exercises
9.1. Prove that $\varepsilon$-approximation of prediction classes implies $\varepsilon T$ regret approximation, as formalized in Lemma 9.2.
9.2. Prove that the class $\prod^{*}$ is optimal for fully observed linear dynamical systems with zero-mean noise.
9.3. Show that the policy class $\prod^{\star}_{h}$ can be learned by policy class $\prod^{L}_{\tilde{h},k,\kappa}$ with $\tilde{h}=0$.
@@
9.5. Consider the square loss in algorithm 6. Give a Newton-method based algorithm based on the Online Newton Step algorithm (see [Haz16]) and write down the resulting regret bound.
9.6. Write down the predictor class $\prod^{\star}$ and $\prod^{\star}_{h}$ for time varying linear dynamical systems. Research the notion of adaptive regret from [Haz16], and reason about conditions under which the class $\prod^{L}$ approximates $\prod^{*}$ for time varying linear dynamical systems.
@@
A key feature of the Kalman filter—and a fundamental limitation—is that it requires knowledge of the system matrices, including both the state transition matrix and the observation model. These matrices encode how the system evolves over time and how observations relate to the hidden state. While the Kalman filter is provably optimal when these parameters are known, real-world applications often involve partially known or entirely unknown system dynamics, making explicit identification challenging.
In contrast, online learning approaches, such as those introduced in Chapter 9, provide a model-free, open-loop alternative by directly learning optimal linear predictors from data. Instead of estimating a latent state and system matrices, these methods construct forecasting models using past observations, making them more adaptive to unknown or time-varying systems. However, these approaches do not incorporate feedback, meaning they operate in an open-loop setting, predicting future states without correcting for new observations.
@@
Beyond classical Kalman filtering, we also introduce a closed-loop online learning method that serves as a convex relaxation of the Kalman filter. This approach retains the feedback mechanism of Kalman filtering while removing the need for explicit system identification, making it applicable in adversarial or uncertain environments. By formulating state estimation as an online convex optimization problem, this method provides provable regret guarantees and serves as a bridge between Bayesian filtering and learning-based estimation techniques.
@@ 10.1 Observable Systems
Observability is a fundamental concept in control theory that determines whether the internal state of a system can be uniquely inferred from its outputs over time. In the infinite-horizon setting, prediction is meaningful only if observations remain bounded, a condition closely related to system controllability (as discussed in Section 4.3).
A general dynamical system maps an initial state to an infinite sequence of observations. If this mapping is injective—meaning that each initial state corresponds to a unique sequence of observations—then the system is said to be observable. Formally:
@@
This system allows only the velocity of the object to be observed, while the position remains hidden. It can be seen that this setting is unobservable, as the initial position cannot be determined based on the velocities and forces applied to the system alone. Indeed, this is a special case of the Galilean Relativity principle of Newtonian mechanics, which states that the laws of physics are the same in all inertial reference frames.
@@
A partially-observed dynamical system specified by system matrcies $(A,B,C)$ is observable if and only if $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$.
@@ 10.2.1 The Power of Linear Predictors
One of the important conclusions from this chapter is that the policy class of linear predictors is optimal with respect to closed loop policies. We prove this by proving two facts:
@@
    The class of Kalman filters is optimal for closed loop policies.
The conclusion from these two facts is that algorithm 6 can be used to learn a policy that competes with the best closed loop optimal predictor in hindsight. If the signal comes from a linear dynamical system subject to Gaussian noise, this means competing with the optimal predictor, which is a Kalman filter. We prove this formally at the end of this section.
10.2.2 Closed Loop Optimal Predictors
@@
So far, we have focused on predicting observations, i.e., $\mathbf{y}_{t}$. However, the Kalman filter also provides optimal state estimates $\hat{\mathbf{x}}_{t}$ given past observations $\mathbf{y}_{t-1:1}$ and controls $\mathbf{u}_{t-1:1}$. This is motivated by the fact that the least-squares optimal predictor of observations can be expressed as:
@@
In Exercise 10.6, we ask the reader to supply a proof of this statement.
@@
Let the control sequence $\mathbf{u}_{t}$ be arbitrary affine functions of $\mathbf{y}_{1:t}$. There exists an efficiently computable sequence of linear predictors $\hat{\mathbf{x}}_{t+1}$ defined recursively as
@@ 10.2.4 Analysis
Since the estimator is recursively defined, a natural strategy is to proceed by induction. Let us say $\hat{\mathbf{x}}_{t}$ is a such a predictor. Further, let
@@
But, when predicting $\mathbf{x}_{t+1}$, we have more information. In particular, we have $\mathbf{y}_{t}$ too. Consider the random variables $\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}$. Can we use this covariate alone to predict $\mathbf{x}_{t+1}$? Note that
@@ 10.2.5 Infinite Horizon Kalman Filtering
However, when does a solution to this Riccati equation exist? A sufficient condition for this is the observability of the underlying dynamical system. In particular, any observable partially-observed linear dynamical system admits a solution to the Riccati equation listed above, and hence, has a steady-state Kalman filter. Mirroring the link between controllability and stabilizability, observability implies that $\rho(A-LC)<1$.
@@ 10.2.6 Linear Predictors and the Kalman Filter
For observable linear dynamical systems, the Kalman filter is contained in the class of linear predictors. This is fomally stated in the following Lemma, whose proof is left as Exercise 10.6.
@@ 10.3 Bayes-Optimality under Gaussian Noise
In this subsection, we further restrict $\mathbf{w}_{t},\mathbf{v}_{t}$ to be i.i.d. draws from Gaussian distributions, and proceed to establish that the Kalman filter in such cases enjoys a stronger notion of optimality; it produces the the best estimator (under a least-squares loss) of the state given past observations among the class of all (not just linear) predictors.  [^1] This is also called the Bayes-optimal predictor. See bibliographic materials for more details and literature. . The setting we consider is that of Gaussian perturbations $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ and known systems matrices $(A,B,C)$. In particular, we consider the case when $\mathbf{w}_{t}\sim\mathcal{N}(0,\Sigma_{x})$ and $\mathbf{v}_{t}\sim\mathcal{N}(0,\Sigma_{y})$ are i.i.d. normally distributed draws, and $x_{0}=0$. Given the realizations $\mathbf{y}_{1},\mathbf{y}_{2}\dots,\mathbf{y}_{t}$ under some control input sequence $\mathbf{u}_{1},\mathbf{u}_{2},\dots\mathbf{u}_{t}$, we would like to compute a prediction $\hat{\mathbf{y}}_{t+1}$ that minimizes the mean square error, which is simply the expected observation as per
@@
for a sequence of matrices $L_{t}\in{\mathbb{R}}^{d_{x}\times d_{y}}$ that can be computed efficiently given $(A,B,C)$, and is independent of the realizations $\mathbf{u}_{1:T},\mathbf{y}_{1:T}$.
@@ 10.4 Conclusion
Beyond its theoretical guarantees, the Kalman filter is widely used in practical applications, ranging from autonomous systems and robotics to finance and neuroscience. However, challenges remain, particularly in scenarios with non-Gaussian noise, time-varying dynamics, or adversarial disturbances. These limitations motivate extensions such as robust filtering, nonlinear filtering (e.g., the extended and unscented Kalman filters), and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
@@ 10.5 Bibliographic Remarks
In 1960, Rudolph Kalman [Kal60] gave a state-space solution to the problem of filtering that was highly amenable to further extensions. See [AM91] for a historical perspective on its development and aftermaths.
@@
  • Nonlinear Filtering Extensions.
    The Kalman filter assumes linear system dynamics and Gaussian noise, limiting its effectiveness in nonlinear settings. The Extended Kalman Filter (see e.g. [UA21]) addresses this by linearizing the dynamics at each time step via a first-order Taylor series expansion. However, the linearization can introduce significant errors in highly nonlinear systems. Notable extensions using more sophisiticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
@@
    For highly nonlinear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative. These methods approximate the posterior distribution using a set of weighted samples (particles) and employ importance sampling and resampling techniques.
@@
    Robust filtering methods seek to provide guarantees under adversarial noise. The $H_{\infty}$ filter [BB08] minimizes the worst-case estimation error rather than the mean-square error, making it suitable for applications with unknown disturbances.
While the Kalman filter remains a cornerstone of state estimation, these alternative methods provide valuable extensions and improvements for nonlinear, adversarial, and data-driven settings. In the next chapter we will study an alternative using a technique which circumvents the need for state estimation altogether, and has stronger guarantees for sequential prediction.

### v5

@@ 9.3.1 Linear Predictors
The simplest prediction rule is a linear function of observations, inputs, or both. More precisely,
@@ 9.3.2 Linear Dynamical Predictors
The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $\mathbf{C}=I$ is the identity matrix. In this situation, the optimal prediction can be calculated using the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$. It can be seen that this prediction rule is optimal for any zero-mean noise or perturbation model. However, this optimality is only true if the state is fully observable. In the next chapter, we consider optimal predictors for zero-mean i.i.d. noise in the more general partially observed case.
@@
Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of states and controls and gives the next output of the system accurately. We denote by $\prod^{\star}$ the set of all such predictors that in the case of zero noise give a perfectly accurate prediction.
@@
With a decaying system, we can limit the memory of the predictor and still maintain accuracy. Thus, it makes sense to consider a more restricted class of predictors that have bounded memory, that is,
@@
where as a matter of convention we say $\mathbf{u}_{t}=0$ for all negative $t$. Denote by $\prod_{h}^{\star}$ the class of all such predictors.
Clearly, for a large $h$ this class of predictors well approximates the true observations $\mathbf{y}_{t}$ if the system is activated without any noise. We next formalize what this means.
@@
$=\left\|\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t-1}\mathbf{x}_{0}-\sum_{i=1:h}CA^{i-1}B\mathbf{u}_{t-i}\right\|$
$=\left\|\sum_{i=h}^{T}CA^{i}B\mathbf{u}_{t-i-1}\right\|$
@@ 9.4 Efficient Learning of Linear Predictors
The previous section gives several natural classes of predictors as well as some of the relationships between them. In particular, the class of linear dynamical predictors is optimal for certain scenarios and requires knowledge of only three matrices. However, the parameterization of this class is nonconvex in the matrices $A,B,C$, which presents a computational challenge.
This challenge is circumvented by using the fact that the class $\prod^{L}$ $\varepsilon$-approximates the class $\prod^{\star}$. In addition, the natural parameterization of $\prod^{L}$ is convex and thus can be learned using online convex optimization.
@@ 9.5 Summary
We began by formulating the problem in terms of partially observable LDS, highlighting the challenges posed by unknown system matrices and noisy observations. We discussed key approaches for online learning in dynamical systems, emphasizing regret minimization as a performance metric. The regret framework allows direct comparisons between learning algorithms by evaluating their cumulative prediction error against the best fixed predictor in hindsight.
@@ 9.6 Bibliographic Remarks
The task of predicting future observations in an online environment is often called “filtering”. Resting at the crossroads between statistical estimation and dynamical systems, the theory of filtering had its early beginnings in the works of Norbert Wiener (see, e.g., [Wie49]).
For a comprehensive treatment of online learning in games and online convex optimization, see [CBL06, Haz16].
In the statistical noise setting, a natural approach for learning is to perform system identification and then use the identified system for prediction. This approach was taken by [SMT+18, SF20, SR19]. The work of [GLS+20] extends identification-based techniques to adversarial noise and marginally stable systems.
The class $\prod^{*}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [AHL+18]. In that paper, it was shown how to learn partially observable and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
Under probabilistic assumptions, the work of [HMR18] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods. Learning without recovery was studied in [GLS+20] even for marginally stable systems. More recently, tensor methods were used in [BLMY23a], and in [BLMY23b] to learn a mixture of linear dynamical systems.
@@ 9.7 Exercises
9.1. Prove that a $\varepsilon$-approximation of prediction classes implies an additional $\varepsilon T$ regret, as formalized in the Lemma 9.2.
9.2. Prove that the class $\prod^{*}$ is optimal for fully observed linear dynamical systems with zero mean noise.
9.3. Show that the policy class $\prod^{\star}_{h}$ can be learned by the policy class $\prod^{L}_{\tilde{h},k,\kappa}$ with $\tilde{h}=0$.
@@
9.5. Consider the square loss in Algorithm 6. Give an algorithm based on the Online Newton Step algorithm (see [Haz16]) and write down the resulting regret bound.
9.6. Write down the predictor class $\prod^{\star}$ and $\prod^{\star}_{h}$ for time-varying linear dynamical systems. Research the notion of adaptive regret from [Haz16], and reason about the conditions under which the class $\prod^{L}$ approximates $\prod^{*}$ for time-varying linear dynamical systems.
@@
A key feature of the Kalman filter and a fundamental limitation is that it requires knowledge of the system matrices, including both the state transition matrix and the observation model. These matrices encode how the system evolves over time and how observations relate to the hidden state. Although the Kalman filter is provably optimal when these parameters are known, real-world applications often involve partially known or entirely unknown system dynamics, making explicit identification challenging.
In contrast, online learning approaches, such as those introduced in Chapter 9, provide a model-free open-loop alternative by directly learning optimal linear predictors from the data. Instead of estimating a latent state and system matrices, these methods construct forecasting models using past observations, making them more adaptive to unknown or time-varying systems. However, these approaches do not incorporate feedback, meaning that they operate in an open-loop setting, predicting future states without correcting for new observations.
@@
Beyond classical Kalman filtering, we also introduce a closed-loop online learning method that serves as a convex relaxation of the Kalman filter. This approach retains the Kalman filtering feedback mechanism, while eliminating the need for explicit system identification, making it applicable in adversarial or uncertain environments. By formulating state estimation as an online convex optimization problem, this method provides provable regret guarantees and serves as a bridge between Bayesian filtering and learning-based estimation techniques.
@@ 10.1 Observable Systems
Observability is a fundamental concept in control theory that determines whether the internal state of a system can be uniquely inferred from its output over time. In the infinite-horizon setting, prediction is meaningful only if observations remain bounded, a condition closely related to system controllability (as discussed in Section 4.3).
A general dynamical system maps an initial state to an infinite sequence of observations. If this mapping is injective, meaning that each initial state corresponds to a unique sequence of observations, then the system is said to be observable. Formally:
@@
This system allows only the velocity of the object to be observed, while the position remains hidden. It can be seen that this setting is unobservable, as the initial position cannot be determined on the basis of the velocities and forces applied to the system alone. Indeed, this is a special case of the Galilean Relativity principle of Newtonian mechanics, which states that the laws of physics are the same in all inertial reference frames.
@@
A partially-observed dynamical system specified by system parameters $(A,B,C)$ is observable if and only if $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$.
@@ 10.2.1 The Power of Linear Predictors
One of the important conclusions from this chapter is that the policy class of linear predictors is optimal with respect to closed-loop policies. We prove this by proving two facts:
@@
    The class of Kalman filters is optimal for closed-loop policies.
The conclusion from these two facts is that algorithm 6 can be used to learn a policy that competes with the best closed-loop optimal predictor in hindsight. If the signal comes from a linear dynamical system subject to Gaussian noise, this means competing with the optimal predictor, which is a Kalman filter. We prove this formally at the end of this section.
10.2.2 Closed-Loop Optimal Predictors
@@
So far, we have focused on predicting observations, that is, $\mathbf{y}_{t}$. However, the Kalman filter also provides optimal state estimates $\hat{\mathbf{x}}_{t}$ given past observations $\mathbf{y}_{t-1:1}$ and controls $\mathbf{u}_{t-1:1}$. This is motivated by the fact that the least-squares optimal predictor of observations can be expressed as:
@@
In Exercise 10.6, we ask the reader to provide a proof of this statement.
@@
Let the control sequence $\mathbf{u}_{t}$ be an arbitrary affine function of $\mathbf{y}_{1:t}$. There exists an efficiently computable sequence of linear predictors $\hat{\mathbf{x}}_{t+1}$ defined recursively as
@@ 10.2.4 Analysis
Since the estimator is defined recursively, a natural strategy is to proceed by induction. Let us say $\hat{\mathbf{x}}_{t}$ is such a predictor. Furthermore, let
@@
But, when predicting $\mathbf{x}_{t+1}$, we have more information. In particular, we also have $\mathbf{y}_{t}$. Consider random variables $\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}$. Can we use this covariate alone to predict $\mathbf{x}_{t+1}$? Note that
@@ 10.2.5 Infinite Horizon Kalman Filtering
However, when does a solution to this Riccati equation exist? A sufficient condition for this is the observability of the underlying dynamical system. In particular, any observable partially-observed linear dynamical system admits a solution to the Riccati equation listed above, and hence, has a steady-state Kalman filter. Mirroring the link between controllability and stabilizability, observability implies $\rho(A-LC)<1$.
@@ 10.2.6 Linear Predictors and the Kalman Filter
For observable linear dynamical systems, the Kalman filter is contained in the class of linear predictors. This is formally stated in the following lemma, whose proof is left as Exercise 10.6.
@@ 10.3 Bayes-Optimality under Gaussian Noise
In this subsection, we further restrict $\mathbf{w}_{t},\mathbf{v}_{t}$ to be i.i.d. draws from Gaussian distributions, and proceed to establish that the Kalman filter in such cases enjoys a stronger notion of optimality; it produces the best estimator (under a least-squares loss) of the state given past observations among the class of all (not just linear) predictors.  [^1] This is also called the Bayes-optimal predictor. See bibliographic materials for more details and literature. . The setting we consider is that of Gaussian perturbations $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ and known systems matrices $(A,B,C)$. In particular, we consider the case where $\mathbf{w}_{t}\sim\mathcal{N}(0,\Sigma_{x})$ and $\mathbf{v}_{t}\sim\mathcal{N}(0,\Sigma_{y})$ are i.i.d. Gaussian and $x_{0}=0$. Given the realizations $\mathbf{y}_{1},\mathbf{y}_{2}\dots,\mathbf{y}_{t}$ under some control input sequence $\mathbf{u}_{1},\mathbf{u}_{2},\dots\mathbf{u}_{t}$, we would like to compute a prediction $\hat{\mathbf{y}}_{t+1}$ that minimizes the mean square error, which is simply the expected observation as per
@@
for a sequence of matrices $L_{t}\in{\mathbb{R}}^{d_{x}\times d_{y}}$ that can be efficiently computed given $(A,B,C)$, and is independent of the realizations $\mathbf{u}_{1:T},\mathbf{y}_{1:T}$.
@@ 10.4 Conclusion
Beyond its theoretical guarantees, the Kalman filter is widely used in practical applications ranging from autonomous systems and robotics to finance and neuroscience. However, challenges remain, particularly in scenarios with non-Gaussian noise, time-varying dynamics, or adversarial disturbances. These limitations motivate extensions such as robust filtering, non-linear filtering (e.g., the extended and unscented Kalman filters) and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
@@ 10.5 Bibliographic Remarks
In 1960, Rudolph Kalman gave a state-space solution to the filtering problem that was highly amenable to further extensions [Kal60]. See [AM91] for a historical perspective on its development and aftermaths.
@@
  • Nonlinear Filtering Extensions. The Kalman filter assumes that the dynamics is linear system and the noise is Gaussian, which limits its effectiveness in nonlinear settings. The Extended Kalman Filter (see, for example, [UA21]) addresses this issue by linearizing the dynamics at each time step via a first-order Taylor series expansion. However, linearization can introduce significant errors in highly non-linear systems. Notable extensions using more sophisticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
@@
    For highly non-linear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative. These methods approximate the posterior distribution using a set of weighted samples (particles) and employ importance sampling and resampling techniques.
@@
    Robust filtering methods seek to provide guarantees under adversarial noise. The $H_{\infty}$ filter [BB08] minimizes the worst-case estimation error rather than the mean-squared error, making it suitable for applications with unknown disturbances.
Although the Kalman filter remains a cornerstone of state estimation, these alternative methods provide valuable extensions and improvements for nonlinear, adversarial, and data-driven settings. In the next chapter, we will study an alternative using a technique which circumvents the need for state estimation altogether and has stronger guarantees for sequential prediction.